A manufacturing production line breaks down two times a week on average. What is the probability that the interarrival times between breakdown will be less than 2 days?
Answer #1


Here, λ = 2 and x = 2
As per Poisson's distribution formula P(X = x) = λ^x * e^(-λ)/x!

We need to calculate P(X <2).
P(X < 2) = (2^0 * e^-2/0!) + (2^1 * e^-2/1!)
P(X < 2) = 0.1353 + 0.2707
P(X < 2) = 0.4060